OpenStudy (anonymous):

PLEASE HELP!!!!!! Given that f(x) = −x + 4 and g(x) = −2x − 3, solve for f(g(x)) when x = 2.

OpenStudy (anonymous):

@tkhunny @ganeshie8 @sidsiddhartha

OpenStudy (anonymous):

f(g(x)) = -g(x) +4 whats g(x)?

OpenStudy (anonymous):

-2x - 3

OpenStudy (anonymous):

put x = 2 in g(x) after putting g(x) in f(x)

OpenStudy (anonymous):

what will be your answer then?

OpenStudy (anonymous):

f(g(x)) = -(g(x)) +4 = - (-2x-3) +4 = 2x +3 +4 = 2x + 7 now put x = 2

OpenStudy (anonymous):

11

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

anytime :)